How do I find the distance between and enter it as a radical or a decimal?
(-5,3) and (5,-3)

Found 2 solutions by josgarithmetic, Theo:
Answer by josgarithmetic(39617)

Look up the Distance Formula. Plug-in the values and compute.

Answer by Theo(13342)

the distance between (-5,3) and (5,-3) would be found by using the formula of:

length of a two dimensional line on a graph = square root of ((y2-y1)^2 + (x2-x1)^2)

you assing (x1,y1) to one of the points and (x2,y2) to the other point.

it doesn't matter which is assigned to which.

the length would be the same.

for example:

let(x1,y1) = (-5,3) and let (x2,y2) = (5,-3).

the length is equal to sqrt((x2-x1)^2 + (y2-y1)^2) = sqrt(5--5)^2 + (-3-3)^2) = sqrt(10^2 + (-6)^2) = sqrt(100 + 36) = sqrt(136).

let (x1,y1) = (5,-3) and let (x2,y2) = (-5,3).

the length is equal to sqrt((x2-x1)^2 + (y2-y1)^2) = sqrt(-5-5)^2 + (3--3)^2) = sqrt((-10)^2 + 6^2) = sqrt(100 + 36) = sqrt(136).

sqrt means the square root of.

if you look at the line on a graph, it would look like this.



the line itself is the hypotenuse of the right triangle formed.

the vertical leg is the length of (y2 - y1).

the horizontal leg is the length of (x2 - x1).

the pythagorean formula is used to get the length of the line itself.